# Robust Distributed Gradient Descent to Corruption over Noisy Channels

### v1

@@ I Introduction
However, there are significant challenges in the use of distributed gradient descent such as the communication overhead [7], the necessity for resilience against system failures [8, 9], and the presence of adversarial attacks [10, 11, 12, 13]. In a distributed framework, workers may become unreliable (called Byzantine workers in the sequel), which can disrupt the learning process. For instance, as demonstrated in [14], the presence of even a single Byzantine worker can undermine the convergence of a distributed gradient descent algorithm.
In this paper, we consider a new problem concerning the distributed gradient descent algorithm but we take into account a novel and critical aspect that has, to the best of our knowledge, not been considered before. We consider the presence of adversarial corruptions at each worker in addition to the usual random (Gaussian) noises in the uplink and downlink channels. The corruptions and noises affect the accuracy of partial gradients and parameters. Each worker experiences adversarial corruption, with the total corruption budget being bounded over the entire time horizon. Our convergence analysis of a modified version of a (lazy) mirror descent-type algorithm addresses the worst-case scenario of these corruptions subject to a long-term constraint on the total corruption budget. We show both theoretically and empirically that our algorithm RDGD and its variants such as RDGD-Restart  can be suitably tuned to be corruption tolerant.
@@ I-A Main Contributions
    We design a mirror descent-inspired distributed gradient descent algorithm RDGD that is robust to corruptions. For smooth and convex functions, we show that the expected suboptimality gap of RDGD  scales as $O\big{(}\frac{C(T)}{m\sqrt{T}}\big{)}$.
@@ I-B Related Works
Another line of research that relevant to our work pertains to Byzantine distributed gradient descent. Byzantine faults refer to the scenario where some nodes in the distributed network act arbitrarily or maliciously, potentially leading to incorrect or suboptimal outcomes. One line of solutions is based on the robust aggregation rules, including trimmed mean [12], coordinate-wise median [12], geometric median [11], and Krum [14]. However, these robust aggregation rules work under the assumption that the number of Byzantine workers is less than half the total number of workers. Our research approach bifurcates from this line of research since we allow the adversary to arbitrarily allocate a bounded amount of corruptions among all workers.
@@ II-B Noisy channel transmission
Now, the noisy partial gradient vector at worker $i$ for based on its local dataset $\mathcal{Z}_{i}$ can be written as
@@ II-D Preliminaries
A differentiable function $f:\Theta\to\mathbb{R}$ is $\mu$-strongly convex on $\Theta$ if for all $\theta,\hat{\theta}\in\Theta$,
@@
Let $\psi:\Theta\to\mathbb{R}$ be a continuously differentiable and strictly convex function. The Bregman divergence $D_{\psi}:\Theta^{2}\rightarrow\mathbb{R}$ is defined as:
@@
The Bregman divergence $D_{\psi}(\theta,\theta^{\prime})$ quantities the gap between $\psi(\theta)$ and its first-order approximation at $\theta^{\prime}$.
@@ III-A The RDGD  Algorithm
At a high level, we utilize a modified lazy mirror descent approach with primal and dual updates. To initialize, we map the initial point $\theta_{0}$ to $z_{0}=\nabla\psi(\theta_{0})$ in the dual space via the mirror map $\nabla\psi$ for some suitably chosen $\mu$-strongly convex function $\psi$. At the $t$-th iteration of RDGD, the parameter server receives the corrupted partial gradients $\tilde{g}_{i,t}$ for all $i\in[m]$. Then, the parameter server calculates the mean of these $\tilde{g}_{i,t}$ and denotes it as $\tilde{g}_{t}=\frac{1}{m}\sum_{i=1}^{m}\tilde{g}_{i,t}$ (Line 11). Next (Line 13), the server updates $z_{0}$ in the dual space to $z_{t}=z_{0}-\sum_{k=1}^{t}\eta_{k}\tilde{g}_{k}.$ It then maps $z_{t}$ back to the point $\theta_{t+1}$ in the primal space (Line 15). The illustration of the primal and dual updates in RDGD is presented in Fig. 2. We let
@@
This process can be considered as a modified version of lazy mirror descent [27]. The intuition behind RDGD  is that in (7), we are minimizing an approximation of the loss function, regularized by the Bregman divergence $D_{\psi}(u,\theta_{0})$. Introducing this regularizer prevents the updated point $\theta_{t+1}$ from straying too far from the current point $\theta_{t}$ when there are corruptions. Also, we denote $\hat{\theta}_{t}=\frac{H_{t-1}}{H_{t}}\hat{\theta}_{t-1}+\frac{\eta_{t}}{H_{t}}\theta_{t},$ (where $H_{t}=\sum_{k=1}^{t}\eta_{k}$) as the output at iteration $t$ (Line 14). We note that $\hat{\theta}_{t}$ is the sum of the weighted gradients based on the history of gradients. As the analysis shows, this ensures that the final output of RDGD $\hat{\theta}_{T}$ is robust against noisy updates.
@@ III-B Convergence Analysis
The promised upper bound is chosen as $U_{t}=L(\hat{\theta}_{t})$, where $\hat{\theta}_{t}$ is a weighted average of the points $(\theta_{k})_{k=1}^{t}$ generated from RDGD  up to $t$-th iteration (specified in Line 14). Given the forms of $U_{t}$ and $L_{t}$, it is evident that $L(\hat{\theta}_{t})-L(\theta^{*})\leqslant G_{t}$. Therefore, demonstrating the convergence of the algorithm primarily is tantamount to establishing a bound for $G_{t}$. We need to track the change of $E_{t}=H_{t}G_{t}-H_{t-1}G_{t-1}$. To facilitate subsequent analyses, we express $G_{t}$ as:
@@
For all $t\geqslant 1$, the following holds:
@@
For all $t\geqslant 1$, the following holds:
$H_{t}G_{t}-H_{t-1}G_{t-1}\leqslant\eta_{t}\langle\tilde{g}_{t}-L(\theta_{t}),\theta^{*}-\theta_{t+1}\rangle\\ -\eta_{t}\langle L(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle-D_{\psi}(\theta_{t+1},\theta_{t}).$
@@
Let $0<M<\infty$ be such that for all $(x,y)\in\mathbb{R}^{p}\times\mathbb{R}$, $\ L(x,y;\cdot)$ is $M$-smooth. Let $\theta_{0}\in\Theta$ be an arbitrary initial point. Let $\{(\theta_{t},z_{t},\hat{\theta}_{t})\}_{t=1}^{T}$ evolve according to RDGD (Algorithm 1) for some $\mu$-strongly convex function $\psi$, where $R_{\Theta}=\max_{\theta\in\Theta}\|\theta-\theta^{*}\|_{2}$. For a fixed number of iterations $T$ and stepsize $\eta_{k}=\frac{1}{\sqrt{T}}$, we have for all $k\geqslant 1$:
$\mathbb{E}[L(\hat{\theta}_{T})-L(\theta^{*})]$ $\leqslant\frac{D_{\psi}(\theta^{*},\theta_{0})}{\sqrt{T}}+\frac{M^{2}}{2\mu}\cdot\frac{1}{\sqrt{T}}$
@@
The first two terms $\frac{D_{\psi}(\theta^{*},\theta_{0})}{\sqrt{T}}$ and $\frac{M^{2}}{2\mu}\cdot\frac{1}{\sqrt{T}}$ are similar to what one would expect for mirror descent [29]. Since we assume that $\Theta$ is bounded, $R_{\Theta}$ is finite. The effect of the adversarial corruption is manifested in the term $\frac{R_{\Theta}C(T)}{m}\cdot\frac{1}{\sqrt{T}}$, where recall that $C(T)$ (defined in (5)) represents the total corruption among all $m$ workers over $T$ iterations. Even though we do not have a lower bound on the gap $\mathbb{E}[L(\hat{\theta}_{T})-L(\theta^{*})]$, the scaling $O\big{(}\frac{C(T)}{m\sqrt{T}}\big{)}$ is intuitively order-optimal as, based on (5), if each worker suffers a constant amount of corruption at each time, then $C(T)=\Theta(m\sqrt{T})$. Besides, if the variances of the noises in both downlink and uplink channels are bounded across all iterations (i.e., $\sigma_{t}^{2}\leqslant\bar{\sigma}^{2}$ for all $t$ and some $\bar{\sigma}^{2}<\infty$), the effect of noisy channels on the bound is the constant $(M+1)\sqrt{p}\sigma R_{\Theta}$. We provide the proof of Theorem 1 in Appendix C.
@@
$\mathbb{E}[L(\hat{\theta}_{t})-L(\theta^{*})]$ $\leqslant\frac{D_{\psi}(\theta^{*},\theta_{0})}{\sqrt{t}}+\frac{M^{2}}{2\mu}\cdot\frac{\log t}{\sqrt{t}}$
@@ IV The Strongly Convex Case
$H_{1}G_{1}\leqslant\frac{\eta_{1}\alpha}{2}\|\theta^{*}-\theta_{1}\|_{2}^{2}+\eta_{1}\langle\tilde{g}_{1}-\nabla L(\theta_{1}),\theta^{*}-\theta_{2}\rangle+\frac{\eta_{1}M^{2}}{4\alpha}.$
@@
Let $\psi(x)=\frac{\eta_{1}\alpha}{2}\|x\|_{2}^{2}$. For all $t\geqslant 1$,
$h_{t}(\theta_{t+1})-h_{t-1}(\theta_{t})\geqslant\eta_{t}\langle\tilde{g}_{t},\theta_{t+1}-\theta_{t}\rangle\\ +\frac{(H_{t}+\eta_{t})\alpha}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}.$
@@
Let $\psi(x)=\frac{\eta_{1}\alpha}{2}\|x\|_{2}^{2}$ and $0\leqslant\frac{\eta_{t}}{H_{t}}\leqslant\frac{\alpha}{M}$. Then, for all $t\geqslant 1$,
$H_{t}G_{t}-H_{t-1}G_{t-1}\leqslant\eta_{t}\langle\tilde{g}_{t}-\nabla L(\theta_{t}),\theta^{*}-\theta_{t+1}\rangle+\frac{\eta_{t}M^{2}}{2\alpha}.$
@@
Let $0<\alpha\leqslant M<\infty$ be such that for all $(x,y)\in\mathbb{R}^{p}\times\mathbb{R}$, $\ L(x,y;\cdot)$ is $M$-smooth and $\alpha$-strongly convex. let $\theta_{0}\in\Theta$ be an arbitrary initial point. Let $\{(\theta_{t},z_{t},\hat{\theta}_{t})\}_{t=1}^{T}$ evolve according to RDGD-SC (Algorithm 2) for $\psi(x)=\frac{\eta_{1}\alpha}{2}\|x\|_{2}^{2}$. Set $R_{\Theta}=\max_{\theta\in\Theta}\|\theta-\theta^{*}\|_{2}$. Let $\eta_{1}=1$ and $\eta_{k}$ satisfy $\eta_{k}\leqslant\frac{\alpha}{M}H_{k}$ for all $k\geqslant 2$. Then for all $t\geqslant 2$:
$\mathbb{E}[L(\hat{\theta}_{t})-L(\theta^{*})]$ $\leqslant\Big{(}\prod_{k=2}^{t}\big{(}1-\frac{\eta_{k}}{H_{k}}\big{)}\Big{)}\frac{\alpha}{2}\|\theta^{*}-\theta_{0}\|_{2}^{2}+\frac{M^{2}}{2\alpha}$
@@
    $\mathbb{E}[L(\hat{\theta}_{t})-L(\theta^{*})]$ $\leqslant\Big{(}1-\frac{\alpha}{M}\Big{)}^{t-1}\frac{\alpha}{2}\|\theta^{*}-\theta_{0}\|_{2}^{2}+\frac{M^{2}}{2\alpha}$
@@
    $\quad+\frac{M^{2}}{2\alpha}+(M+1)\sqrt{p}\sigma R_{\Theta}.$ (14)

### v2

@@ I Introduction
However, there are significant challenges in the use of distributed gradient descent such as the communication overhead [7], the necessity of having resilience against system failures [8, 9], and the presence of adversarial attacks [10, 11, 12, 13]. In a distributed framework, workers may become unreliable (called Byzantine workers in the sequel), which can disrupt the learning process. For instance, as demonstrated in [14], the presence of even a single Byzantine worker can undermine the convergence of a distributed gradient descent algorithm.
In this paper, we consider a new problem concerning the distributed gradient descent algorithm but we take into account a novel and critical aspect that has, to the best of our knowledge, not been considered before. We consider the presence of adversarial corruptions at each worker in addition to the usual random (Gaussian) noises in the uplink and downlink channels. The corruptions and noises affect the accuracy of partial gradients and parameters. Each worker experiences adversarial corruption, with the total corruption budget being bounded over the entire time horizon. An application of our work could be in a wireless setting with analog transmission of the gradients. However, instead of assuming that this wireless setting is corruption-free, we assume that there are malicious parties in the system that systematically and deterministically modify the gradients that are sent to a central server for downstream machine learning tasks. Our convergence analysis of a modified version of a (lazy) mirror descent-type algorithm addresses the worst-case scenario of these corruptions subject to a long-term constraint on the total corruption budget. We show both theoretically and empirically that our algorithm RDGD and its variants such as RDGD-Restart  can be suitably tuned to be corruption tolerant.
@@ I-A Main Contributions
    We design a mirror descent-inspired distributed gradient descent algorithm RDGD that is robust to corruptions. For smooth and convex functions, we show that the expected suboptimality gap of RDGD  decays as $O\big{(}\frac{1}{\sqrt{T}}+\frac{C(T)}{m\sqrt{T}}\big{)}$.
@@ I-B Related Works
Another line of research that relevant to our work pertains to Byzantine distributed gradient descent. Byzantine faults refer to the scenario where some nodes in the distributed network act arbitrarily or maliciously, potentially leading to incorrect or suboptimal outcomes. One line of solutions is based on the robust aggregation rules, including trimmed mean [12], coordinate-wise median [12], geometric median [11], and Krum [14]. However, these robust aggregation rules work under the assumption that the number of Byzantine workers is less than half the total number of workers. Our research approach bifurcates from this line of research since we allow the adversary to arbitrarily allocate a bounded amount of corruptions among all workers. Our model, however, naturally precludes the possibility of tolerating arbitrary perturbations as median-based approaches that eliminate corruptions of arbitrarily large magnitudes cannot be directly employed.
@@ II-B Noisy channel transmission
Now, the noisy partial gradient vector at worker $i$ based on its local dataset $\mathcal{Z}_{i}$ can be written as
@@ II-D Preliminaries
Definition 2 (Lipschitz Continuity)
A function $f:\Theta\to\mathbb{R}$ is $K$-Lipschitz continuous on $\Theta$ if for all $\theta,\hat{\theta}\in\Theta$,
$\|f(\hat{\theta})-f(\theta)\|_{2}\leqslant K\|\hat{\theta}-\theta\|_{2}.$
@@
A function $f:\Theta\to\mathbb{R}$ is $\mu$-strongly convex on $\Theta$ if for all $\theta,\hat{\theta}\in\Theta$,
@@
Let $\psi:\Theta\to\mathbb{R}$ be a continuously differentiable and strictly convex function called the mirror map. The Bregman divergence $D_{\psi}:\Theta^{2}\rightarrow\mathbb{R}$ is defined as:
@@
The Bregman divergence $D_{\psi}(\theta,\theta^{\prime})$ quantities the gap between $\psi(\theta)$ and its first-order Taylor approximation at $\theta^{\prime}$.
@@ III-A The RDGD  Algorithm
At a high level, we utilize a modified lazy mirror descent approach with primal and dual updates. To initialize, we map the initial point $\theta_{0}$ to $z_{0}=\nabla\psi(\theta_{0})$ in the dual space via the mirror map $\nabla\psi$ for some suitably chosen $\mu$-strongly convex function $\psi$. At the $t$-th iteration of RDGD, the parameter server receives the corrupted partial gradients $\tilde{g}_{i,t}$ for all $i\in[m]$. Then, the parameter server calculates the mean of these $\tilde{g}_{i,t}$ and denotes it as $\tilde{g}_{t}=\frac{1}{m}\sum_{i=1}^{m}\tilde{g}_{i,t}$ (Line 11).
Next (Line 13), the server updates $z_{0}$ in the dual space to $z_{t}=z_{0}-\sum_{k=1}^{t}\eta_{k}\tilde{g}_{k}.$ It then maps $z_{t}$ back to the point $\theta_{t+1}$ in the primal space (Line 15). The illustration of the primal and dual updates in RDGD is presented in Fig. 2. We let
@@
This process can be considered as a modified version of lazy mirror descent [27]. The intuition behind RDGD  is that in (7), we are minimizing an approximation of the loss function, regularized by the Bregman divergence $D_{\psi}(u,\theta_{0})$. The mirror descent procedure operates in a dual space where the effects of corrupted gradients can be modulated differently than if done in the primal space. This reduces the impact of the adversarial corruptions since the updates are not directly in the parameter space where the adversary might have more influence. Regularization makes the optimization process more sensitive to changes in directions that are more likely to be influenced by Byzantine workers, effectively dampening their impact. By shaping the update dynamics through regularization, we limit the influence of corrupted gradients without significantly hindering the convergence speed of legitimate updates.
We denote $\hat{\theta}_{t}=\frac{H_{t-1}}{H_{t}}\hat{\theta}_{t-1}+\frac{\eta_{t}}{H_{t}}\theta_{t},$ (where $H_{t}=\sum_{k=1}^{t}\eta_{k}$) as the output at iteration $t$ (Line 14). We note that $\hat{\theta}_{t}$ is the sum of the weighted gradients based on the history of the gradients. As the analysis shows, this ensures that the final output of RDGD $\hat{\theta}_{T}$ is robust against noisy updates.
The best choice of the mirror map $\psi$ depends on the specifics of the task. Commonly used mirror maps include:
  1. 1.
    $\ell_{2}$-norm squared $\psi(x)\,\propto\,\|x\|_{2}^{2}$: Suitable for general optimization due to the balance between computational efficiency and being amenable to deriving convergence guarantees.
  2. 2.
    $\ell_{1}$-norm $\psi(x)=\|x\|_{1}$: Promotes sparsity in solutions, making it advantageous in high-dimensional settings.
  3. 3.
    Negative Entropy $\psi(x)=\sum_{i}x_{i}\log x_{i}$: Effective for tasks involving probability distributions or simplex constraints.
Choosing the appropriate mirror map impacts the algorithm’s performance and robustness by tailoring the geometry of the optimization landscape to the task’s requirements. In Section VI-A, we experimentally compare the efficacy of different mirror maps for our specific setting and conclude that the $\ell_{2}$-norm squared is the most appropriate, at least for the least squares regression task.
@@ III-B Convergence Analysis
The promised upper bound is chosen as $U_{t}=L(\hat{\theta}_{t})$, where $\hat{\theta}_{t}$ is a weighted average of the points $(\theta_{k})_{k=1}^{t}$ generated from RDGD  up to $t$-th iteration (specified in Line 14). Given the forms of $U_{t}$ and $L_{t}$, it is evident that $L(\hat{\theta}_{t})-L(\theta^{*})\leqslant G_{t}$. Therefore, demonstrating the convergence of the algorithm is tantamount to establishing a bound for $G_{t}$. We need to track the change of $E_{t}=H_{t}G_{t}-H_{t-1}G_{t-1}$. To facilitate subsequent analyses, we express $G_{t}$ as:
@@
For all $t\geqslant 2$, the following holds:
@@
For all $t\geqslant 2$, the following holds:
$H_{t}G_{t}-H_{t-1}G_{t-1}\leqslant\eta_{t}\langle\tilde{g}_{t}-\nabla L(\theta_{t}),\theta^{*}-\theta_{t+1}\rangle\\ -\eta_{t}\langle\nabla L(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle-D_{\psi}(\theta_{t+1},\theta_{t}).$
@@
Let $0<K,M<\infty$ be such that for all $(x,y)\in\mathbb{R}^{p}\times\mathbb{R}$, $\ L(x,y;\cdot)$ are $M$-smooth and $K$-Lipschitz continuous. Let $\theta_{0}\in\Theta$ be an arbitrary initial point. Let $\{(\theta_{t},z_{t},\hat{\theta}_{t})\}_{t=1}^{T}$ evolve according to RDGD (Algorithm 1) for some $\mu$-strongly convex function $\psi$, where $R_{\Theta}=\max_{\theta\in\Theta}\|\theta-\theta^{*}\|_{2}$. Let $\sigma_{k}\leqslant\sigma$ for all $k\geqslant 1$. For a fixed number of iterations $T$ and stepsize $\eta_{k}=\frac{1}{\sqrt{T}}$, we have for all $k\geqslant 1$:
$\mathbb{E}[L(\hat{\theta}_{T})-L(\theta^{*})]$ $\leqslant\frac{D_{\psi}(\theta^{*},\theta_{0})}{\sqrt{T}}+\frac{K^{2}}{2\mu}\cdot\frac{1}{\sqrt{T}}$
@@
The first two terms $\frac{D_{\psi}(\theta^{*},\theta_{0})}{\sqrt{T}}$ and $\frac{K^{2}}{2\mu}\cdot\frac{1}{\sqrt{T}}$ are similar to what one would expect for mirror descent [29]. Since we assume that $\Theta$ is bounded, $R_{\Theta}$ is finite. The effect of the adversarial corruption is manifested in the term $\frac{R_{\Theta}C(T)}{m}\cdot\frac{1}{\sqrt{T}}$, where recall that $C(T)$ (defined in (5)) represents the total corruption among all $m$ workers over $T$ iterations. Even though we do not have a lower bound on the gap $\mathbb{E}[L(\hat{\theta}_{T})-L(\theta^{*})]$, the scaling $O\big{(}\frac{C(T)}{m\sqrt{T}}\big{)}$ is intuitively order-optimal as, based on (5), if each worker suffers a constant amount of corruption at each time, then $C(T)=\Theta(m\sqrt{T})$. Besides, if the variances of the noises in both downlink and uplink channels are bounded across all iterations (i.e., $\sigma_{t}^{2}\leqslant\bar{\sigma}^{2}$ for all $t$ and some $\bar{\sigma}^{2}<\infty$), the effect of noisy channels on the bound is $(M+1)\sqrt{p}\sigma R_{\Theta}$. The proof of Theorem 1 is provided in Appendix C.
@@
$\mathbb{E}[L(\hat{\theta}_{t})-L(\theta^{*})]$ $\leqslant\frac{D_{\psi}(\theta^{*},\theta_{0})}{\sqrt{t}}+\frac{K^{2}}{2\mu}\cdot\frac{\log t}{\sqrt{t}}$
@@ IV The Strongly Convex Case
$H_{1}G_{1}\leqslant\frac{\eta_{1}\alpha}{2}\|\theta^{*}-\theta_{1}\|_{2}^{2}+\eta_{1}\langle\tilde{g}_{1}-\nabla L(\theta_{1}),\theta^{*}-\theta_{2}\rangle+\frac{\eta_{1}K^{2}}{4\alpha}.$
@@
Let $\psi(x)=\frac{\eta_{1}\alpha}{2}\|x\|_{2}^{2}$. For all $t\geqslant 2$,
$h_{t}(\theta_{t+1})-h_{t-1}(\theta_{t})$ $\geqslant\eta_{t}\langle\tilde{g}_{t},\theta_{t+1}-\theta_{t}\rangle$ (12)
$\qquad+\frac{H_{t}\alpha}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}.$ (13)
@@
Let $\psi(x)=\frac{\eta_{1}\alpha}{2}\|x\|_{2}^{2}$ Then, for all $t\geqslant 2$,
$H_{t}G_{t}-H_{t-1}G_{t-1}\leqslant\eta_{t}\langle\tilde{g}_{t}-\nabla L(\theta_{t}),\theta^{*}-\theta_{t+1}\rangle+\frac{\eta_{t}K^{2}}{2\alpha}.$
@@
Let $0<\alpha\leqslant M<\infty$ and $0<K<\infty$ be such that for all $(x,y)\in\mathbb{R}^{p}\times\mathbb{R}$, $\ L(x,y;\cdot)$ is $M$-smooth, $K$-Lipschitz continuous and $\alpha$-strongly convex. let $\theta_{0}\in\Theta$ be an arbitrary initial point. Let $\{(\theta_{t},z_{t},\hat{\theta}_{t})\}_{t=1}^{T}$ evolve according to RDGD-SC (Algorithm 2) for $\psi(x)=\frac{\eta_{1}\alpha}{2}\|x\|_{2}^{2}$. Let $R_{\Theta}=\max_{\theta\in\Theta}\|\theta-\theta^{*}\|_{2}$ and $\sigma_{k}\leqslant\sigma$ for all $k\geqslant 1$. Let $\eta_{1}=1$ and $\eta_{k}$ satisfy $\eta_{k}\leqslant\frac{\alpha}{M}H_{k}$ for all $k\geqslant 2$. Then for all $t\geqslant 2$:
$\mathbb{E}[L(\hat{\theta}_{t})-L(\theta^{*})]$ $\leqslant\Big{(}\prod_{k=2}^{t}\big{(}1-\frac{\eta_{k}}{H_{k}}\big{)}\Big{)}\frac{\alpha}{2}\|\theta^{*}-\theta_{0}\|_{2}^{2}+\frac{K^{2}}{2\alpha}$
@@
    $\mathbb{E}[L(\hat{\theta}_{t})-L(\theta^{*})]$ $\leqslant\Big{(}1-\frac{\alpha}{M}\Big{)}^{t-1}\frac{\alpha}{2}\|\theta^{*}-\theta_{0}\|_{2}^{2}+\frac{K^{2}}{2\alpha}$
@@
    $\quad+\frac{K^{2}}{2\alpha}+(M+1)\sqrt{p}\sigma R_{\Theta}.$ (16)